# subfigure0

### v1

@@ 5 Discussion
    Other complexity measures. This work focused on Krylov complexity, but there are other definitions of complexity, including out-of-time-order correlation functions which are characterized by the Lyapunov exponent. It would be interesting to see how the Lyapunov exponent responds to supersymmetry and its breaking in $\mathcal{N}=2$ SYK. For example, without supersymmetry, the mass deformation causes the Lyapunov exponent to go to zero at low temperatures, while the Krylov exponent continues to be near-maximal [Chapman_2024]. It would be interesting to see if this behavior persists in the supersymmetric case. Other complexity measures, including circuit complexity and Nielsen complexity, can also be explored [Baiguera_2025, Balasubramanian2021, Craps2024].
@@
    Random matrix theory. The results of this paper could be compared with predictions from random matrix theory [Kanazawa_2017, Kar_2022]. This could provide further insights into the universality classes of quantum chaos and complexity in deformed $\mathcal{N}=2$ SYK.
@@
    Symmetry-resolved Krylov complexity. Another way to study the effect of symmetries on Krylov complexity is to analyze symmetry-resolved Krylov complexity, which captures the impact each symmetry sector has on complexity [Caputa_2025, Caputa_2025_2]. Generally, complexity at early times equals the average over symmetry sectors, but at late times the interplay between sectors becomes important. Symmetry-resolved Krylov complexity of $\mathcal{N}=2$ SYK can be studied across sectors of fixed $U(1)$ or $U(1)_{R}$ charge. Breaking supersymmetry usually also breaks these symmetries, but it would be interesting to study the contribution of each sector to the overall complexity when supersymmetry is intact.
@@
    Holographic dual. This paper studied the quantum mechanical notion of complexity, but the most interesting area of future research would be to explore the holographic analogue of Krylov complexity in the dual gravitational system [Jian_2021, Rabinovici_2023, Ambrosini_2025, Ambrosini_2025_2, Caputa_2024]. In a two-sided system consisting of two copies of the SYK model, the holographic dual to Krylov complexity has been interpreted as the size of the wormhole connecting the boundaries. The supersymmetric version of this picture involves super-JT gravity and the size of supersymmetric wormholes. It would be interesting to interpret the deformations of supersymmetric SYK treated here in terms of the growth of supersymmetric wormholes in the bulk.
@@
Data access statement. The code and data used in this study are publicly available on the arXiv page of this paper.
@@ Appendix B Comparison of Different Lanczos Algorithms
To reduce the computational cost of computing the Krylov basis, the partial reorthogonalization algorithm reorthogonalizes the new basis element only when the level of overlap with previous basis elements exceeds a certain threshold. To achieve this, PRO estimates the overlap between the $n$th Krylov basis element and the $k$th element using the recurrence relation [Rabinovici_2021]
@@
Figure 22 compares the FO and PRO algorithms for the computation of the Lanczos coefficients of the $\mathcal{N}=2$ SYK model, with and without deformation. The parameters of the PRO algorithm were set to $\epsilon_{\text{P}}=10^{-13}$ and $\epsilon_{\text{M}}=10^{-15}$, and $n_{\text{O}}=2$. The left column shows the absolute percent error in the Lanczos coefficients $b_{n}$ between the two algorithms, while the right column shows the cumulative number of reorthogonalizations performed by the PRO algorithm. When $\epsilon=0$, the percent error is extremely small, on the order of $10^{-14}$, until around $n\sim 100$. After $n\sim 100$, the error grows to around 1%, while the number of reorthogonalizations increases. After 1000 iterations, the PRO algorithm has only performed around 100 reorthogonalizations, while keeping the error to order 1%.
When the UV deformation is turned on, the PRO algorithm is able to keep the error within around 1% using only 10 reorthogonalizations, two orders of magnitude fewer than the FO algorithm. When the mass deformation is used, PRO tends to keep the error around $10^{-14}$, but reorthogonalizes as much as in the undeformed case. The error is also consistently low for the $\mathcal{N}=0$ model, with order 100 reorthogonalizations in 1000 iterations. These experiments demonstrate the effectiveness of the PRO algorithm in computing the Lanczos coefficients of the $\mathcal{N}=2$ SYK model using significantly fewer reorthogonalizations than the FO algorithm. The FO algorithm took roughly 700700800800 s to compute $1000$ iterations on a MacBook Pro with M1 Pro chip and $32$  GB RAM, while PRO took 500500600600 s. While this is a modest speedup, it grows significantly as the number of iterations increases. Use of the PRO algorithm enabled the computation of the full Lanczos sequences for $N=8$ reported in the main text.
Figure 22: Percent error in the Lanczos coefficients between the FO and PRO Lanczos algorithms $E_{b}$ and the cumulative number of PRO reorthogonalizations $N_{r}$ as the Lanczos iteration $n$ increases. The parameters of the PRO algorithm were set to $\epsilon_{\text{P}}=10^{-13}$ and $\epsilon_{\text{M}}=10^{-15}$. The number of full orthogonalizations used in both algorithms was set to $n_{\text{O}}=2$. The values $N=8$, $q=3$, and $J=1$ were used throughout.
@@ Appendix C Estimation of Krylov Complexity Parameters
Figure 24 shows the estimates of $\alpha(t)$ given by Equation (C.4). The values of $\alpha$ at early times correspond to the growth rate in the quadratic phase, while the values at late times correspond to the growth rate in the linear phase, denoted $\upsilon_{K}$ in the main text. Because these direct estimates can be noisy, an alternative method to estimate $\alpha$ and $\upsilon_{K}$ was employed based on $\delta(t)$. In this method, the times at which $\delta(t)$ crosses the thresholds of $0.5$ and 0 were identified. In practice, the thresholds were adjusted to account for noise in the estimates. The two crossing times define the intervals over which quadratic and linear fits to the complexity data were performed. To improve robustness, multiple fits were carried out by varying the interval endpoints, and the final estimates were taken as the averages over all fits. The results are summarized in Figure 12 of the main text.

### v2

@@ 5 Discussion
    Other complexity measures. This work focused on Krylov complexity, but there are other definitions of complexity, including out-of-time-order correlation functions which are characterized by the Lyapunov exponent. It would be interesting to see how the Lyapunov exponent responds to supersymmetry and its breaking in $\mathcal{N}=2$ SYK. For example, without supersymmetry, the mass deformation causes the Lyapunov exponent to go to zero at low temperatures, while the Krylov exponent continues to be near-maximal [20]. It would be interesting to see if this behavior persists in the supersymmetric case. Other complexity measures, including circuit complexity and Nielsen complexity, can also be explored [10, 11, 23].
@@
    Random matrix theory. The results of this paper could be compared with predictions from random matrix theory [41, 42]. This could provide further insights into the universality classes of quantum chaos and complexity in deformed $\mathcal{N}=2$ SYK.
@@
    Symmetry-resolved Krylov complexity. Another way to study the effect of symmetries on Krylov complexity is to analyze symmetry-resolved Krylov complexity, which captures the impact each symmetry sector has on complexity [16, 17]. Generally, complexity at early times equals the average over symmetry sectors, but at late times the interplay between sectors becomes important. Symmetry-resolved Krylov complexity of $\mathcal{N}=2$ SYK can be studied across sectors of fixed $U(1)$ or $U(1)_{R}$ charge. Breaking supersymmetry usually also breaks these symmetries, but it would be interesting to study the contribution of each sector to the overall complexity when supersymmetry is intact.
@@
    Holographic dual. This paper studied the quantum mechanical notion of complexity, but the most interesting area of future research would be to explore the holographic analogue of Krylov complexity in the dual gravitational system [40, 59, 6, 7, 15, 34, 33, 3, 1, 5, 2]. In a two-sided system consisting of two copies of the SYK model, the holographic dual to Krylov complexity has been interpreted as the size of the wormhole connecting the boundaries. The supersymmetric version of this picture involves super-JT gravity and the size of supersymmetric wormholes [4]. It would be interesting to interpret the deformations of supersymmetric SYK treated here in terms of the growth of supersymmetric wormholes in the bulk.
@@
Data access statement. The code and data used in this study are publicly available at https://arxiv.org/abs/2511.20769.
@@ Appendix B Comparison of Different Lanczos Algorithms
To reduce the computational cost of computing the Krylov basis, the partial reorthogonalization algorithm reorthogonalizes the new basis element only when the level of overlap with previous basis elements exceeds a certain threshold. To achieve this, PRO estimates the overlap between the $n$th Krylov basis element and the $k$th element using the recurrence relation [56]
@@
[…]
Regardless of the Lanczos algorithm used, computing the Lanczos coefficients often requires high numerical precision to maintain stability. Insufficient precision can cause quantities such as $a_{n}\equiv(\mathcal{O}_{n}|\mathcal{L}|\mathcal{O}_{n})$, which should vanish for Hermitian operators $\mathcal{O}_{n}$, to drift away from zero. This drift produces inaccurate Lanczos coefficients, particularly after many iterations. Low precision can also lead to the total number of iterations exceeding the theoretical maximum Krylov dimension reported in Figure 6. This effect is illustrated in Figure 23, which compares the Lanczos coefficients computed using the FO algorithm with two different levels of numerical precision for the UV-deformed $\mathcal{N}=2$ SYK model with $\epsilon=0.6$. The high-precision computation terminates within the Krylov bound of $241$, while the low-precision computation exceeds the bound. This example highlights the importance of using sufficient numerical precision when computing the Lanczos coefficients to ensure accurate results. Utilizing high precision also increases the runtime of the algorithms, effectively limiting the system sizes that can be simulated.
Figure 22: The cumulative number of PRO reorthogonalizations $N_{r}$ as the Lanczos iteration $n$ increases for $\mathcal{N}=2$ SYK. The parameters were set to $\epsilon_{\text{P}}=10^{-13}$, $\epsilon_{\text{M}}=10^{-15}$, and $n_{\text{O}}=2$. The first value of $\epsilon$ in the legend corresponds to the UV deformation, and the second value to the mass deformation. The values $N=6$, $q=3$, and $J=1$ were used throughout.
Figure 23: Comparison of the Lanczos coefficients $b_{n}$ computed using the FO algorithm using two different levels of numerical precision for the UV-deformed $\mathcal{N}=2$ SYK model with $\epsilon=0.6$. The low-precision computation used the IEEE 754 double-precision floating-point format, which utilizes between $15$ and $16$ decimal places, or dps. The high-precision computation used $1024$ dps. The values $N=6$, $q=3$, and $J=1$ were used throughout.
@@ Appendix C Estimation of Krylov Complexity Parameters
Figure 25 shows the estimates of $\alpha(t)$ given by Equation (C.4). The values of $\alpha$ at early times correspond to the growth rate in the quadratic phase, while the values at late times correspond to the growth rate in the linear phase, denoted $\upsilon_{K}$ in the main text. Because these direct estimates can be noisy, an alternative method to estimate $\alpha$ and $\upsilon_{K}$ was employed based on $\delta(t)$. In this method, the times at which $\delta(t)$ crosses the thresholds of $0.5$ and zero were identified. In practice, the thresholds were adjusted to account for noise in the estimates. The two crossing times define the intervals over which quadratic and linear fits to the complexity data were performed. To improve robustness, multiple fits were carried out by varying the interval endpoints, and the final estimates were taken as the averages over all fits. The results are summarized in Figure 12 of the main text.